# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v2

@@ 4.2 Generalization to Multi-Machine Case
  • External Delay: The external delay for any large job $j$ is bounded by $O(\sqrt{n/m})\cdot p_{j}$. In the single-machine setting, each job has $\sqrt{n}$ insertion choices, resulting in an expected delay affecting $\sqrt{n}$ jobs. In the multi-machine setting, we allocate $\sqrt{n/m}$ choices to $w_{j}$ and $m$ choices to $m_{j}$, resulting in an expected delay affecting only $\sqrt{n/m}$ jobs.
@@ 5 Deterministic Non-Preemptive Algorithm
This section introduces a deterministic scheduling algorithm that operates without preemption or restart. In the single-machine setting, the trivial $O(n)$-competitive algorithm is known to be tight. However, when the number of machines $m$ is no longer constant, this $O(n)$ bound becomes loose. Our algorithm achieves a competitive ratio of $O(n/m^{2}+\sqrt{n/m}\log m)$.
The algorithm builds on the online rank-based partitioning framework introduced in Algorithm 1. Upon the arrival of a job $j$, it is classified as either large or small according to the rule specified in Algorithm 1 with $\ell=\lfloor\sqrt{nm}\rfloor$. Recall that this partitioning framework not only distinguishes between large and small jobs but may also release proxy jobs and mark certain jobs as proxied.
Machines are assigned specific roles as follows:
  • Machine 1: Reserved for large jobs.
  • Mixed machines (Machines $2$ to $\lceil m/2\rceil+1$): Accept both small and large jobs.
  • Small-Only machines (Machines $\lceil m/2\rceil+2$ to $m$): Restricted to small jobs only.
Note that not all Mixed machines are available to large jobs at all times. Instead, we adopt a fine-grained policy to dynamically determine the number of active Mixed machines dedicated to large jobs. This is generated by a function $\gamma$, which depends on $k$, the number of active large jobs still waiting. Roughly speaking, at any given time, we activate $\gamma(k)$ Mixed machines to serve large jobs, where:
$\gamma(k)=\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor.$
Finally, the scheduling policy follows a generalized version of NSJF, respecting the machine roles defined above:
  • If Machine 1 is idle, it schedules the smallest waiting large job, if available.
  • If a Mixed machine is idle, it first schedules the smallest waiting small job, if one is available. Otherwise, let $\theta$ be the number of Mixed machines currently processing large jobs. If $\theta<\gamma(k)$, where $k$ is the number of active large jobs still waiting, the algorithm schedules the smallest such job.
  • If a Small-Only machine is idle, it schedules the smallest waiting small job, if available.
@@
$\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
@@
We first bound the waiting time component $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if proxied) of the flow time which does not include the $p_{j}$ part. At any time $t$ in this interval, $j$ is not scheduled because all active Mixed machines are busy. Let $\theta(t)$ be the number of Mixed machines currently processing jobs, and let $k(t)$ be the number of active large jobs waiting at time $t$. By the scheduling rule, we either have
$\gamma(k(t))\leq\theta(t)+1,\quad\text{or}\quad\theta(t)=\lceil m/2\rceil.$
In the second case, since $k(t)$ is at most $\lfloor\sqrt{nm}\rfloor$, and using the definition $\gamma(k)=\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor$, we get:
$k(t)\leq O(\sqrt{n/m})\cdot(\theta(t)+1).$
Also, note that while a large job is waiting, the large-only machine (Machine 1) must be busy. Letting $\mathbb{I}_{k(t)\geq 1}(t)$ denote the indicator function for whether $k(t)\geq 1$, we have:
@@
Combining all these elements and reincorporating $p_{j}$, we have:
@@
$\leq\sum_{j\in L}p_{j}+O(\sqrt{n/m})\cdot\sum_{j\in J}p_{j}$
$=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
Next, we analyze the contribution of small jobs. The first loss occurs because of the large-only machine. Ideally, even without any large job, our algorithm is a (m1)(m-1)-machine NSJF algorithm, which is competitive to the (m1)(m-1)-machine preemptive optimal solution. However, the (m1)(m-1)-machine preemptive optimal may already be larger than the $m$-machine optimal solution. To address it, we give the following lemma.
@@
For every $0\leq k\leq m-1$, letting $\textup{{OPT}}_{m-k}$ be the optimal preemptive solution only using mkm-k machines, we have $\textup{{OPT}}_{m-k}=O\left(\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}$.
Before proceeding, we recall that the preemptive optimal solutions discussed here, $\textup{{OPT}}_{m}$ and $\textup{{OPT}}_{m-k}$, operate under preemptive model without migration.
@@
Consider a preemptive optimal solution with $m$ machines. We will use it to construct a feasible solution using mkm-k machines and prove that its total flow time increases by at most $\frac{nk}{m(m-k)}\cdot\textup{{OPT}}_{m}$. This will complete the proof of the lemma.
First, we sort the machines in the optimal solution $\textup{{OPT}}_{m}$ by their total workloads. Let $J_{i}$ denote the set of jobs scheduled on machine $i$ in $\textup{{OPT}}_{m}$, and let $J_{[i]}=\bigcup_{h=1}^{i}J_{h}$ denote the union of jobs on the first $i$ machines. Machines are sorted so that
$\sum_{j\in J_{1}}p_{j}\leq\sum_{j\in J_{2}}p_{j}\leq\cdots\leq\sum_{j\in J_{m}}p_{j}.$
We consider the first $k$ machines and reassign all jobs in $J_{[k]}$ to the remaining mkm-k machines in arbitrary order. By construction, we have:
$\sum_{j\in J_{[k]}}p_{j}\leq\frac{k\sum_{j\in J}p_{j}}{m}.$
The reassignment proceeds as follows. For each job $j\in J_{[k]}$, we preserve its starting time $t$ from $\textup{{OPT}}_{m}$, and insert it at the same time $t$ on one of the machines among machines mk+1m-k+1 to $m$. As a result, the existing jobs on that machine are delayed by at most $p_{j}$.
To minimize the impact, we always place $j$ on the machine with the fewest currently assigned jobs. Since there are mkm-k machines and at most $n$ jobs, this ensures the number of jobs on that machine is at most n/(mk)n/(m-k). Therefore, the additional cost due to delaying existing jobs is at most $np_{j}/(m-k)$.
On the other hand, since we insert job $j$ at its original start time, its flow time is at most what it was in $\textup{{OPT}}_{m}$. Even though future delays may increase $j$’s flow time, these increases are charged to the delay costs of subsequent jobs. Thus, the increase in total flow time due to inserting a single job $j$ compared to $\textup{{OPT}}_{m}$ is at most
$\frac{np_{j}}{m-k}.$
Summing over all $j\in J_{[k]}$, the total increase in flow time is at most
$\frac{n\sum_{j\in J_{[k]}}p_{j}}{m-k}\leq\frac{nk\sum_{j\in J}p_{j}}{m(m-k)}\leq\frac{nk\cdot\textup{{OPT}}_{m}}{m(m-k)}.$
In conclusion, we obtain the desired bound:
$\textup{{OPT}}_{m-k}\leq\left(1+\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}.$
As in Lemma 12, we illustrate the ideas again here for completeness. We partition the timeline into intervals based on when large jobs are started. Let $t_{h}$ denote the start time of the $h$-th large job, with $t_{0}=0$. Let $S_{h}$ be the set of small jobs released during the interval $t_{h-1},t_{h})$.
Let $\textup{{OPT}}_{m-1}(S_{h})$ denote the optimal total flow time for scheduling $S_{h}$ on m1m-1 machines. Since the intervals are disjoint, we have:
$\textup{{OPT}}_{m-1}\geq\sum_{h}F(\textup{{OPT}}_{m-1}(S_{h})).$
We compare our algorithm’s flow time on $S_{h}$ to $\textup{{OPT}}_{m-1}(S_{h})$. During $t_{h-1},t_{h})$, no new large jobs are started, and all jobs in $S_{h}$ arrive after $t_{h-1}$. The only difference from running $\textup{{NSJF}}(S_{h})$ on a fresh instance is that each machine may still be processing a job that started before $t_{h-1}$. We refer to these as initial blocking jobs, and define the corresponding initial blocking period on machine $i$ as $t_{h-1},t_{h-1}+b_{i})$, where $b_{i}$, the initial blocking time is the remaining processing time at $t_{h-1}$. Thus, our algorithm behaves like NSJF on m1m-1 machines with such initial blocking times periods.
This behavior is captured by Lemma 29. To apply it, we define a shifted instance $S^{\prime}_{h}$ by translating the time origin to $t_{h-1}$, i.e., replacing each release time $r_{j}$ with $r_{j}-t_{h-1}$. Then, the flow time of each job $j\in S_{h}$ under our algorithm matches its flow time in $\textup{{NSJF}}(S^{\prime}_{h},\vec{b})$, where each machine has initial blocking interval $0,b_{i})$.
Lemma 10.
For each small-job period $(t_{h-1},t_{h}]$, the total length of initial blocking time satisfies:
$\sum_{i=1}^{m-1}b_{i}\leq O\left(\sqrt{\frac{m}{n}}\log m\right)\cdot\textup{{OPT}}.$
@@
We split the analysis based on whether the blocking job is small or large.
We first consider the blocking caused by small jobs. By Lemma 1, each small job has size at most $O\left(\frac{\textup{{OPT}}}{\sqrt{nm}}\right)$. Since at most $m$ machines can be blocked, the total blocking time from small jobs is:
$O\left(\frac{\textup{{OPT}}}{\sqrt{nm}}\right)\cdot m=O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right).$
For the blocking caused by large jobs, we proceed as follows. For large jobs causing blocking, let $p_{i}$ be the size of the $i$-th large job that causes blocking. When it starts, at least i1i-1 Mixed machines are already processing large jobs, so it must satisfy $\gamma(k)\geq i$ for the number of active large jobs still waiting $k$. Using $\gamma(k)=\lfloor k/(\sqrt{n/m})\rfloor$, we get $k\geq i\sqrt{n/m}$, and thus:
$p_{i}\leq\frac{\textup{{OPT}}}{k}\leq\frac{\textup{{OPT}}}{i\sqrt{n/m}}\leq\frac{1}{i}\cdot O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right).$
Summing over $i=1$ to $m$, the total blocking from large jobs is:
$\sum_{i=1}^{m}p_{i}\leq O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right)\cdot\sum_{i=1}^{m}\frac{1}{i}=O\left(\sqrt{\frac{m}{n}}\log m\cdot\textup{{OPT}}\right).$
Adding both parts completes the proof. ∎
With the bound on the initial blocking times $b_{i}$, we can apply the generalized NSJF analysis in Lemma 29. This lemma allows us to compare the performance of NSJF with initial blocking times against the optimal schedule without any blocking. We now prove the main result.
@@
There exists an online polynomial time deterministic non-preemptive algorithm for online total flow time minimization with competitive ratio $O(n/m^{2}+\sqrt{n/m}\log m)$, against the preemptive offline optimal solution.
@@
Let $\tau$ be the size upper bound for small jobs, i.e., $\tau=O(\textup{{OPT}}/\sqrt{nm})$. For each small-job segment $t_{h-1},t_{h})$, our algorithm performs the same as $\textup{{NSJF}}(S^{\prime}_{h})$ with initial blocking vector $\vec{b}$, denoted by $F_{j}(\textup{{NSJF}}(S^{\prime}_{h},\vec{b}))$. By Lemma 29, and letting $B=\sum_{i}b_{i}$, we have:
$\sum_{j\in S^{\prime}_{h}}F_{j}(\textup{{NSJF}}(S^{\prime}_{h},\vec{b}))\leq\textup{{OPT}}_{m-1}(S_{h})+2|S_{h}|\tau+\frac{|S_{h}|B}{m}.$
@@
$\sum_{j\in S}F_{j}\leq F(\textup{{OPT}}_{m-1})+2n\tau+\frac{nB}{m}.$
Using the bounds $\tau=O(\textup{{OPT}}/\sqrt{nm})$ and $B=O(\sqrt{m/n}\log m\cdot\textup{{OPT}})$, we get:
@@ 6 Deterministic Algorithm with Kill-and-Restart
To utilize the power of kill-and-restart without relying on randomization, we first introduce a fine-grained online partitioning framework with the following modification (see Algorithm 3 for details, and note that the threshold will be set to $\sqrt{nm}$):
  • Modified partition rule: Let $P$ be the total size of all previously released jobs (excluding proxy jobs), which serves as a lower bound on OPT. We classify a new job $j$ as a small job if $p_{j}\leq 4P/\ell$, rather than using the large job queue as in the original rule.
The modification is motivated by the need for a size-based lower bound in the restart setting, as opposed to a purely rank-based one. For instance, we aim to avoid killing jobs that are too small. These changes are mild and preserve all the original properties of the online partitioning framework.
@@
Function Partition($i,\ell$):
 // New Rule: An absolute smallness rule to filter out tiny jobs.
@@
    return $i$ is small
 else if $|A|<\ell$ then
      Add $i$ to $A$;
    return $i$ is large
 else if $p_{i}>\min_{j\in A}p_{j}$ then
@@
      Remove $j$ from $A$ and add $i$ to $A$;
@@
       // The scheduler will actually process $j$ when it processes $j^{\prime}$.
@@
    return $i$ is large
 else
    return $i$ is small
@@
Under the online rank-based partitioning framework (fine-grained for restart), the following properties hold:
@@
    The number of large jobs is at most $\ell$ at any time.
@@
    Let $S$ be the set of small jobs, including proxy jobs. Then $S$ is strictly weaker than the full job set $J$, meaning $F(\textup{{OPT}}(S))\leq F(\textup{{OPT}}(J))$.
@@
    The size of any small job, including any retired large job, is upper bounded by $O(\textup{{OPT}}/\ell)$.
@@
    New size-based lower bound: The size of any large job is at least $4P(r_{j})/\ell$, where $P(r_{j})$ denotes the total processing time of all jobs released up to $r_{j}$, excluding proxy jobs.
@@
After introducing the fine-grained partitioning framework, we now describe the basic design of our scheduling policy, which divides the $m$ machines into two roles:
  • Small-Only machines (Machines $1$ to $\lfloor m/2\rfloor$): These machines accept only small jobs.
  • Mixed machines (Machines $\lfloor m/2\rfloor+1$ to $m$): These machines can process both small and large jobs.
Our scheduling policy is a generalized version of the NSJF algorithm, incorporating the specific machine assignment rule described above. In particular, the algorithm makes scheduling decisions based on the partitioning produced by Algorithm 1. Each job is classified as either large or small, and further as retired or active, and proxied or unproxied.
The scheduling rule proceeds as follows: the algorithm always selects the smallest unproxied waiting job to schedule. Additionally, Small-Only machines are restricted to processing only small jobs.
The kill-and-restart rule applies only to active large jobs. Intuitively, an active large job should be killed only if it is blocking too many small jobs. We say that an active large job blocks a small job if it is processing at the small job’s release time. To detect such situations, the algorithm monitors the number of blocked small jobs. Specifically, if at any point the number of blocked small jobs reaches the threshold $\sqrt{nm}$, we kill all currently processing active large jobs. At the moment of a kill event, as well as when a new large job starts processing, we reset the counter tracking the number of blocked small jobs.
With this, we complete the high-level description of our algorithmic design. The full pseudocode is presented in Algorithm 4.
@@
Data: Potential counter $\phi$ initiated as 0; every job’s current states as determined by Algorithm 3.
@@
    $\phi\leftarrow\phi+1$ if $i$ is blocked ;
    if $\text{$\phi$}=\lfloor\sqrt{nm}\rfloor$ then
@@
         Schedule $j$ on $i$ ;
       $\phi\leftarrow 0$ if $j$ is large ;
@@
Moving into the analysis, we now distinguish between several types of job sets: The real job set $J$, which includes only actual jobs and excludes proxy jobs. The unblocked small job set $S$, and the blocked small job set $B$, both of which may include proxy jobs. The large job set $L$, which includes all large jobs, even if they become retired or are proxied at some point.
Note that we do not have $B\cup S\cup L\subseteq J$ due to the presence of proxy jobs. We analyze the total flow time contributed by $S$, $B$, and $L$ separately.

### v3

@@ 4.2 Generalization to Multi-Machine Case
  • External Delay: The external delay for any large job $j$ is bounded by $O(\sqrt{n/m})\cdot p_{j}$. In the single-machine setting, each job has $\sqrt{n}$ insertion choices, resulting in an expected delay affecting $\sqrt{n}$ jobs. In the multi-machine setting, we allocate $\sqrt{n/m}$ choices to $w_{j}$ and $m$ choices to $m_{j}$, resulting in an expected delay affecting only $O(\sqrt{n/m})$ jobs.
@@ 5 Deterministic Non-Preemptive Algorithm
This section presents a deterministic non-preemptive algorithm with a competitive ratio of $O(n/m^{2}+\sqrt{n/m}\log m)$. This nearly closes the gap between the trivial $O(n/m)$ upper bound and the $\Omega(n/m^{2})$ lower bound of [10], particularly when $m$ is large relative to $n$.
The algorithm builds on the online rank-based partitioning framework (Algorithm 1) with $\ell=\lfloor\sqrt{nm}\rfloor$. Based on this classification, machines are assigned different roles to handle large and small jobs separately, as described below.
The first machine, called the Large-Only machine, is dedicated exclusively to large jobs and schedules the smallest waiting large job if available. For the remaining Mixed machines $2\dots m$, a dynamic threshold function
$\gamma(k)=\min\left\{\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor,\lceil m/2\rceil\right\}.$
controls how many of them may process large jobs, where $k$ is the number of active large jobs currently waiting. When one of these Mixed machines becomes idle, it schedules the smallest waiting job regardless of its size if the number of Mixed machines currently processing large jobs is less than $\gamma(k)$; otherwise, it schedules small jobs only using NSJF. Since there are at most $\lceil m/2\rceil$ Mixed machines running large jobs at any time, there are always $m-1-\lceil m/2\rceil$ machines exclusively dedicated to small jobs.
@@
$\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n/m})\cdot\textup{{OPT}}.$
@@
Consider the waiting time of a large job $j$ during the interval $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if proxied). At any time $t$ in this interval, $j$ is an active large job that has not been scheduled, either because all Mixed machines are busy, or because at least $\gamma(k(t))$ of them are already processing large jobs.
Let $\theta(t)$ denote the number of busy Mixed machines at time $t$ (processing any job, large or small), and let $k(t)$ denote the number of active large jobs waiting at time $t$. Since at least $\gamma(k(t))$ of these machines are processing large jobs, we have by the scheduling rule:
$\theta(t)=m-1\quad\text{or}\quad\gamma(k(t))\leq\theta(t).$
By the definition of $\gamma$, this gives $k(t)=O(\sqrt{n/m})\cdot(\theta(t)+1)$.
Whenever $k(t)\geq 1$, the Large-Only machine must also be busy, so the total number of busy machines is at least $\theta(t)+1$. Since each busy machine contributes to the total processing time:
@@
Combining with the processing time of the jobs themselves:
@@
$\leq\sum_{j\in L}p_{j}+O(\sqrt{n/m})\cdot\sum_{j\in J}p_{j}=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
Next, we analyze the contribution of small jobs. Since the Large-Only machine is reserved for large jobs, small jobs are effectively scheduled on m1m-1 machines. To compare against the $m$-machine optimum, we first bound the cost of losing one machine.
@@
For every $0\leq k\leq m-1$, the optimal preemptive flow time on mkm-k machines satisfies $\textup{{OPT}}_{m-k}\leq\left(1+\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}$, where both optima are under the preemptive model without migration.
@@
Starting from an optimal preemptive schedule on $m$ machines, we construct a feasible schedule on mkm-k machines. Sort the machines by total workload and let $J_{[k]}$ denote the jobs on the $k$ lightest machines. Their total processing time satisfies $\sum_{j\in J_{[k]}}p_{j}\leq k\sum_{j\in J}p_{j}/m$.
Each job $j\in J_{[k]}$ is reassigned to one of the remaining mkm-k machines, inserted at its original start time. This delays existing jobs on the target machine by at most $p_{j}$. To minimize impact, each job is placed on the machine with the fewest currently assigned jobs, ensuring at most n/(mk)n/(m-k) jobs per machine. The additional flow time from inserting $j$ is therefore at most $np_{j}/(m-k)$.
Summing over all reassigned jobs:
$\textup{{OPT}}_{m-k}\leq\textup{{OPT}}_{m}+\frac{n\sum_{j\in J_{[k]}}p_{j}}{m-k}\leq\textup{{OPT}}_{m}+\frac{nk}{m(m-k)}\cdot\textup{{OPT}}_{m}.$
Our key tool for analyzing small jobs is the generalized NSJF analysis (Lemma 29, proved in Appendix A), which extends Lemma 2 to handle initial blocking periods. If NSJF runs on $m^{\prime}$ machines where machine $i$ is unavailable during $0,b_{i})$, and all job sizes are bounded by $\tau$, then
$F(\textup{{NSJF}}(\vec{b}))\leq F(\textup{{OPT}}(\vec{0}))+2n\tau+\frac{nB}{m^{\prime}},$
where $B=\sum_{i}b_{i}$ is the total blocking time and $\textup{{OPT}}(\vec{0})$ is the optimal preemptive schedule on $m^{\prime}$ machines without blocking.
To apply this, we partition the timeline into intervals $t_{h-1},t_{h})$, where $t_{h}$ is the start time of the $h$-th large job on an Mixed machine (excluding the Large-Only machine) and $t_{0}=0$. Let $S_{h}$ denote the small jobs released during $t_{h-1},t_{h})$. Since no Mixed machine starts a new large job during $t_{h-1},t_{h})$, all m1m-1 Mixed machines are available for small jobs, and every job in $S_{h}$ is scheduled before $t_{h}$. Some machines may be initially blocked by jobs still in progress from the previous interval: machine $i$ is unavailable during $t_{h-1},t_{h-1}+b_{h,i})$, where $b_{h,i}$ is its remaining processing time at $t_{h-1}$. Shifting the time origin to $t_{h-1}$, the flow time of $S_{h}$ under our algorithm matches that of $\textup{{NSJF}}(S^{\prime}_{h},\vec{b_{h}})$, where $S^{\prime}_{h}$ is the shifted instance and $\vec{b_{h}}$ is the blocking vector.
It remains to bound the total blocking time.
Lemma 10 (Blocking Time).
For each interval $(t_{h-1},t_{h}]$, the total initial blocking time satisfies:
$\sum_{i=1}^{m-1}b_{i}=O\left(\sqrt{\frac{m}{n}}\log m\right)\cdot\textup{{OPT}}.$
@@
At time $t_{h-1}$, each Mixed machine is blocked by at most one job still in progress. We bound the total blocking separately for small and large blocking jobs.
Small blocking jobs.
Each small job has size at most $O(\textup{{OPT}}/\sqrt{nm})$ by Lemma 1. Since at most m1m-1 machines can be blocked, the total blocking from small jobs is at most $O(\sqrt{m/n}\cdot\textup{{OPT}})$.
Large blocking jobs.
Order the large jobs that are still running on Mixed machines at time $t_{h-1}$ by their start times, and let $p_{i}$ denote the size of the $i$-th such job. When the $i$-th job was started, at least i1i-1 other Mixed machines were already processing large jobs, so the scheduling rule required $\gamma(k)\geq i$, where $k$ is the number of active large jobs waiting at that moment. By the definition of $\gamma$, this implies $k\geq i\sqrt{n/m}$. Since these $k$ waiting jobs are all at least as large as $p_{i}$ (otherwise they would have been scheduled first), we have $k\cdot p_{i}\leq\textup{{OPT}}$, giving:
$p_{i}\leq\frac{\textup{{OPT}}}{i\sqrt{n/m}}=\frac{1}{i}\cdot O\left(\sqrt{\frac{m}{n}}\right)\cdot\textup{{OPT}}.$
There are at most $\lceil m/2\rceil$ such jobs, so the total blocking from large jobs is:
$\sum_{i=1}^{\lceil m/2\rceil}p_{i}\leq O\left(\sqrt{\frac{m}{n}}\right)\cdot\textup{{OPT}}\cdot\sum_{i=1}^{\lceil m/2\rceil}\frac{1}{i}=O\left(\sqrt{\frac{m}{n}}\log m\right)\cdot\textup{{OPT}}.$
Combining both parts gives the claimed bound. ∎
With the blocking time bounded, Lemma 29 applies directly. We now prove the main result.
@@
There exists an online polynomial-time deterministic non-preemptive algorithm for total flow time minimization that is $O(n/m^{2}+\sqrt{n/m}\log m)$-competitive against the preemptive offline solution.
@@
Let $\tau$ be the size upper bound for small jobs, i.e., $\tau=O(\textup{{OPT}}/\sqrt{nm})$. For each small-job segment $t_{h-1},t_{h})$, our algorithm performs the same as $\textup{{NSJF}}(S^{\prime}_{h})$ with initial blocking vector $\vec{b_{h}}$. By Lemma 29 on m1m-1 machines, let $B_{h}=\sum_{i}b_{h,i}$:
$\sum_{j\in S_{h}}F_{j}\leq F(\textup{{OPT}}_{m-1}(S_{h}))+2|S_{h}|\tau+\frac{|S_{h}|B_{h}}{m-1}.$
@@
$\sum_{j\in S}F_{j}\leq F(\textup{{OPT}}_{m-1})+2n\tau+\frac{n\cdot\max_{h}B_{h}}{m-1}.$
Using the bounds $\tau=O(\textup{{OPT}}/\sqrt{nm})$ and $\max_{h}B_{h}=O(\sqrt{m/n}\log m\cdot\textup{{OPT}})$, we get:
@@ 6 Deterministic Algorithm with Kill-and-Restart
To utilize the power of kill-and-restart without relying on randomization, we first introduce a refined online partitioning framework with the following modification (see Algorithm 3 for details, and note that the threshold will be set to $\sqrt{nm}$):
Modified partition rule.
Let $P$ be the total size of all previously released jobs (excluding proxy jobs), which serves as a lower bound on OPT. We classify a new job $j$ as a small job if $p_{j}\leq 4P/\ell$, and large otherwise. The other procedures remain identical to Algorithm 1.
This modification preserves the original properties of the framework. The additional size filter provides a lower bound on the processing time of large jobs, which will be used in the analysis of kill events (1).
@@
Procedure Partition($i,\ell$):
@@
      Mark $i$ as small;
 else
    if $|A|=\ell$ then
@@
       // $j$ will be retired.
         Remove $j$ from AA;
@@
          // The scheduler will actually process $j$ when processing $j^{\prime}$.
@@
      end if
     Add $i$ to AA;
      Mark $i$ as large;
@@
Under the online rank-based partitioning framework (refined for restart), the following properties hold:
@@
    The number of active large jobs (the set $A$) is at most $\ell$ at any time.
@@
    The processing time of any job classified as small, or any retired large job, is upper bounded by $O(\textup{{OPT}}/\ell)$.
@@
    Let $S$ be the set of all jobs ever classified as small, including all proxy jobs. Let $J$ be the original job set. The instance $S$ is no harder to schedule than $J$, meaning $F(\textup{{OPT}}(S))\leq F(\textup{{OPT}}(J))$.
@@
    The processing time of any large job $j$ is at least $4P(r_{j})/\ell$, where $P(r_{j})$ denotes the total processing time of all jobs released up to $r_{j}$, excluding proxy jobs.
@@
After introducing the refined partitioning framework, we now describe the basic design of our scheduling policy, which divides the $m$ machines into two roles:
  • Small-Only machines ($1$ to $\lfloor m/2\rfloor$): These machines accept only small jobs.
  • Mixed machines ($\lfloor m/2\rfloor+1$ to $m$): These machines accept both small and large jobs.
Our scheduling policy builds on the partitioning produced by Algorithm 3. Each job is classified as either large or small; each large job is further classified as active or retired, and as proxied or unproxied.
Scheduling rule.
When an Mixed machine becomes idle, it schedules the smallest waiting unproxied job (regardless of whether it is large or small). When a Small-Only machine becomes idle, it schedules the smallest waiting small job.
Kill-and-restart rule.
We say that a newly released small job is blocked if any active large job is currently being processed, and unblocked otherwise. The algorithm maintains a counter $\phi$, which is incremented by one for each blocked small job and reset to zero whenever an active large job begins processing or is killed. When $\phi$ reaches $\lfloor\sqrt{nm}\rfloor$, the algorithm kills all currently processing active large jobs and resets $\phi$ to zero.
The full pseudocode is presented in Algorithm 4.
@@
Data: Blocked small job counter $\phi$ initialized to 0; every job’s current states as determined by Algorithm 3.
@@
 if $j\neq\bot$, i.e., there exists a new small job $j$ then
[…]
    $\phi\leftarrow\phi+1$ if $j$ is blocked ;
    if $\phi=\lfloor\sqrt{nm}\rfloor$ then
@@
         Schedule $j$ on machine $i$ ;
       $\phi\leftarrow 0$ if $j$ is an active large job ;
@@
The basic invariant we aim to keep in the algorithm is Lemma 12.
Lemma 12.
The number of waiting blocked small jobs at any time is at most $\sqrt{nm}$.
Proof.
We partition the timeline into intervals, where each interval begins when the algorithm starts processing an active large job and ends just before the next such event. By the algorithm, the counter $\phi$ is reset to 0 at the start of each interval. Within an interval, every newly released blocked small job increments $\phi$ by one, and when $\phi$ reaches $\lfloor\sqrt{nm}\rfloor$, all currently processing active large jobs are killed and $\phi$ is reset. Therefore, at most $\lfloor\sqrt{nm}\rfloor$ blocked small jobs are created in each interval.
It remains to show that all blocked small jobs created in one interval have started processing before the next interval begins. After a kill event (or if the active large jobs complete naturally), no active large job is being processed. At this point, every waiting blocked small job is smaller than every waiting active large job, so by the shortest-first scheduling rule, all blocked small jobs are scheduled before any active large job can start processing again. ∎
Moving into the analysis, we now distinguish between several types of job sets: The real job set $J$, which includes only actual jobs and excludes proxy jobs. The unblocked small job set $S$, and the blocked small job set $K$, both of which may include proxy jobs. The large job set $L$, which includes all large jobs, even if they become retired at some point.
Note that $K\cup S\cup L\not\subseteq J$ with the presence of proxy jobs. We analyze the total flow time contributed by $S$, $K$, and $L$ separately.